# VPTracker: Global Vision-Language Tracking via Visual Prompt and MLLM

### v1

@@ Abstract
Vision-Language Tracking aims to continuously localize objects described by a visual template and a language description. Existing methods, however, are typically limited to local search, making them prone to failures under viewpoint changes, occlusions, and rapid target movements. In this work, we introduce the first global tracking framework based on Multimodal Large Language Models (VPTracker), exploiting their powerful semantic reasoning to locate targets across the entire image space. While global search improves robustness and reduces drift, it also introduces distractions from visually or semantically similar objects. To address this, we propose a location-aware visual prompting mechanism that incorporates spatial priors into the MLLM. Specifically, we construct a region-level prompt based on the target’s previous location, enabling the model to prioritize region-level recognition and resort to global inference only when necessary. This design retains the advantages of global tracking while effectively suppressing interference from distracting visual content. Extensive experiments show that our approach significantly enhances tracking stability and target disambiguation under challenging scenarios, opening a new avenue for integrating MLLMs into visual tracking. Code is available at https://github.com/jcwang0602/VPTracker.
I Introduction
@@
Visual object tracking (VOT) is a fundamental problem in computer vision, supporting a wide range of downstream applications such as intelligent surveillance, human–robot interaction, augmented reality, and autonomous systems [10149530]. Traditional visual trackers primarily rely on an initial visual template and perform frame-by-frame localization using appearance similarity [ye2022ostrack]. Although such methods have achieved strong short-term performance, they are inherently constrained by the limited representational capacity of a single visual template. As a result, their robustness deteriorates in complex scenarios involving drastic appearance changes, heavy occlusions, large camera motion, or long-term absence.
@@
To overcome the ambiguity of purely visual cues, vision-language tracking has recently emerged as a promising direction [11209477]. By incorporating natural language descriptions, trackers can leverage high-level semantic priors to better distinguish objects with similar appearance and maintain identity consistency over time. This paradigm has shown particular advantages in multi-object scenes and cases where visual cues alone are insufficient [wang2021towards]. However, existing most vision-language trackers still predominantly adopt local search strategies, restricting the search region to a small neighborhood around the previous target location. While efficient, this localized assumption makes the tracker vulnerable to drift: once the target moves rapidly, undergoes abrupt viewpoint changes, or becomes temporarily occluded, the tracker may fail to rediscover it.
@@
Recent advancements in Multimodal Large Language Models (MLLMs) have opened new possibilities for rethinking the tracking pipeline. MLLMs exhibit remarkable capabilities in global scene understanding, cross-modal reasoning [bai2025qwen3]. These strengths suggest that MLLMs could fundamentally reshape long-term tracking, allowing the model to perform global search over the entire image rather than relying on local heuristics. In this paper, we present the first global vision–language tracking framework powered by Multimodal Large Language Models (MLLMs), termed VPTracker. Leveraging the strong semantic reasoning capabilities of MLLMs, our approach can localize the target across the entire image space. Despite their impressive semantic understanding, we observe that large-scale language models inherently lack temporal modeling and spatial continuity awareness—two key elements required for reliable visual tracking. To overcome this limitation, we introduce a location-aware visual prompting mechanism that injects spatial priors directly into the MLLM’s inference process. Specifically, we construct a region-level prompt based on the target’s previous location, guiding the model to focus on the most probable search area while still maintaining the ability to perform global reasoning when necessary. This design preserves the robustness of global search while substantially reducing false positives caused by visually or semantically similar distractors.
@@
The experimental results on the commonly used visual language tracking datasets TNL2K [wang2021towards], and TNLLT [wang2025reasoningtrack] demonstrate that our method has achieved significant improvements in robustness, target re-identification, and long-term stability. More importantly, VPTracker introduces a new perspective of integrating MLLM into vision-language tracking tasks, indicating that future tracking will not only be driven by appearance similarity but also by advanced cross-modal reasoning. In summary, the main contributions of this work are threefold:
@@
  • Extensive experiments on TNL2K and TNLLT demonstrate that our approach achieves superior robustness, improved recognition capability, and strong long-term stability, establishing a new direction for integrating MLLMs into vision–language tracking..
II Related Work
II-A Object Tracking
Visual object tracking (VOT) has been extensively studied over the past decade, with research evolving from correlation filter–based methods to modern deep learning–based frameworks [ye2022ostrack]. Early trackers such as KCF [henriques2014high] and ECO [danelljan2017eco] model target appearance through handcrafted features and efficient frequency-domain operations. With the success of deep convolutional networks, Siamese-based trackers (e.g., SiamFC [bertinetto2016fully], SiamRPN [li2018high], SiamMask [hu2023siammask]) became dominant due to their strong template-matching capability and real-time performance. Transformer-based trackers (e.g., TransT [chen2021TransT], STARK [tang2022SFTransT], OSTrack [ye2022ostrack]) further improve robustness by modeling long-range dependencies with self-attention mechanisms.
@@ II-A Object Tracking
II-B Multimodal Large Language Models
Multimodal Large Language Models (MLLMs), such as GPT-4V [achiam2023gpt], LLaVA [li2024llava], Qwen-VL[bai2025qwen3], and InternVL [wang2025internvl3], have exhibited strong capabilities in high-level scene understanding, cross-modal reasoning, and open-vocabulary recognition [wang2025unlocking]. Unlike traditional vision-language models that output fixed embeddings, MLLMs are able to perform explicit reasoning, generate spatial descriptions, and interpret complex relationships between objects. These properties make them promising for challenging video tasks such as grounding, captioning, and instruction-following.
@@ III-A Problem Formulation
III-B Global Tracking via MLLMs
To address the aforementioned issues, this paper proposes a global tracking method based on MLLM. Leveraging MLLM’s powerful reasoning capabilities, it searches for targets across the entire video frame given a visual template and corresponding linguistic description, and provides bounding boxes for the detected targets. As shown in Fig. 1, our method relies solely on a single multimodal large language model and does not require any additional components.
@@ III-B Global Tracking via MLLMs
III-C Visual Prompt based on Location-aware
@@
IV Experiment
IV-A Implementation Details
@@
IV-B Comparison with State-of-the-art Methods
@@
The TNLLT dataset is a recently proposed benchmark for long-term vision–language tracking, designed to assess a model’s capability to localize target objects over extended video sequences using natural language descriptions. Its test set consists of 50 long and diverse video sequences. To better assess robustness, TNLLT further provides 15 challenging attributes, such as occlusion, deformation, motion blur, illumination variation, fast motion, low resolution, out-of-view, and background clutter. These rich annotations make TNLLT a comprehensive benchmark for evaluating long-term temporal modeling, and cross-modal reasoning in vision–language tracking. We compare VPTracker with existing state-of-the-art trackers in Tab.I. Compared to the state-of-the-art ReasoningTrack, which uses MLLM to update the linguistic description of the object during tracking, we achieve significant improvements over prior methods.
@@
TABLE I: Overall tracking performance on TNLLT dataset. The best results are highlighted in bold.
@@
OSTrack [ye2022ostrack] ECCV2022 $B$ 52.1 57.3 63.6
MixFormer [cui2022mixformer] CVPR2022 $B$ 56.0 61.1 67.8
AiATrack [gao2022aiatrack] ECCV2022 $B$ 53.0 56.9 63.5
JointNLT [zhou2023joint] CVPR2023 NL 45.0 47.9 55.5
JointNLT [zhou2023joint] CVPR2023 BL 45.4 48.2 55.9
All-in-one [zhang2023all] MM2023 BL 49.7 54.8 60.6
MMTrack [zheng2023toward] TCSVT2023 BL 55.8 61.8 67.7
CiteTrack [citetracker] ICCV2023 $B$ 50.3 55.3 60.8
ROMTrack [cai2023robust] ICCV2023 $B$ 49.0 54.8 59.9
GRM [gao2023generalized] CVPR2023 $B$ 50.0 55.4 61.3
ODTrack [zheng2024odtrack] AAAI2024 $B$ 54.4 61.3 65.9
EVPTrack [shi2024evptrack] AAAI2024 $B$ 54.4 61.3 65.8
UVLTrack [ma2024unifying] AAAI2024 $B$ 53.1 58.1 64.1
AQATrack [xie2024autoregressive] CVPR2024 $B$ 55.1 61.5 66.5
UVLTrack [ma2024unifying] AAAI2024 NL 51.9 57.4 63.1
UVLTrack [ma2024unifying] AAAI2024 BL 54.4 60.3 66.2
SUTrack [sutrack] AAAI2025 BL 50.8 55.4 60.0
LMTrack [xu2025less] AAAI2025 $B$ 48.7 53.0 59.8
CTVLT [feng2025enhancing] ICASSP2025 BL 60.9 69.5 73.6
@@
The TNL2K dataset is a large dataset used for training and evaluating Vision-Language tracking. It contains 2000 video sequences and 663 words, of which 1300 videos are used for training and 700 videos are used for testing. It introduces two new challenges: adversarial samples and modality switches. We compare VPTracker with existing advanced trackers in Tab.II. we achieve significant improvements over prior methods.
TABLE II: Overall tracking performance on TNL2K dataset. The best results are highlighted in bold.
@@
TrDiMP [wang2021transformer] CVPR2021 $B$ - - -
TransT [tang2022SFTransT] CVPR2021 $B$ 50.7 51.7 57.1
AutoMatch [zhang2021learn] AAAI2021 $B$ 47.2 43.5 -
SNLT [feng2021siamese] CVPR2021 BL 27.6 41.9 -
CapsuleTNL [ma2021capsule] MM 2021 BL - - -
VLTTT [guo2022divert] NeurIPS2022 BL 53.1 53.3 -
CTRNLT [li2022cross] CVPR2022 BL 44.0 45.0 52.0
JointNLT [zhou2023joint] CVPR2023 BL 56.9 58.1 73.6
OneTrack [hong2024onetracker] CVPR2024 BL 58.0 59.1 -
QueryNLT [shao2024context] CVPR2024 BL 57.8 58.7 75.6
MAVLT [shi2025mamba] TCSVT2025 BL 63.1 66.7 -
ProVLT [10947500] TCSVT2025 BL 59.8 61.8 -
TemTrack [xie2025robust] AAAI 2025 $B$ 58.8 - -
@@ IV-C Qualitative Comparison
IV-D Ablation study
To validate the effectiveness of our proposed location-aware visual prompting mechanism, we conduct comparisons on the TNL2K dataset with and without the visual prompt. As shown in Tab. III, incorporating the visual prompting mechanism leads to consistent and notable improvements across metrics such as AUC, PR, and NPR. To more clearly demonstrate the effectiveness of our proposed VP mechanism, we present a visual comparison of the tracking results in Fig. 4. Since the textual description of the target may only be applicable to the first frame, global tracking without the position-aware visual prompt can lead to incorrect prediction when visually similar objects appear in the scene.
TABLE III: Overall tracking performance on TNL2K dataset.

### v2

@@ Abstract
Vision-Language Tracking aims to continuously localize objects described by a visual template and a language description. Existing methods, however, are typically limited to local search, making them prone to failures under viewpoint changes, occlusions, and rapid target movements. In this work, we introduce the first global tracking framework based on Multimodal Large Language Models (VPTracker), exploiting their powerful semantic reasoning to locate targets across the entire image space. While global search improves robustness and reduces drift, it also introduces distractions from visually or semantically similar objects. To address this, we propose a location-aware visual prompting mechanism that incorporates spatial priors into the MLLM. Specifically, we construct a region-level prompt based on the target’s previous location, enabling the model to prioritize region-level recognition and resort to global inference only when necessary. This design retains the advantages of global tracking while effectively suppressing interference from distracting visual content. Extensive experiments show that our approach significantly enhances tracking stability and target disambiguation under challenging scenarios, opening a new avenue for integrating MLLMs into visual tracking. Code and models are available at https://github.com/jcwang0602/VPTracker.
keywords:
[…]
1 Introduction
@@
Visual object tracking (VOT) is a fundamental problem in computer vision, supporting a wide range of downstream applications such as intelligent surveillance, human–robot interaction, augmented reality, and autonomous systems 10149530 ; ZHAO202679 ; GE202615 ; XING202559 . Traditional visual trackers primarily rely on an initial visual template and perform frame-by-frame localization using appearance similarity ye2022ostrack . Although such methods have achieved strong short-term performance, they are inherently constrained by the limited representational capacity of a single visual template. As a result, their robustness deteriorates in complex scenarios involving drastic appearance changes, heavy occlusions, large camera motion, or long-term absence.
@@
To overcome the ambiguity of purely visual cues, vision-language tracking has recently emerged as a promising direction 11209477 . By incorporating natural language descriptions, trackers can leverage high-level semantic priors to better distinguish objects with similar appearance and maintain identity consistency over time. This paradigm has shown particular advantages in multi-object scenes and cases where visual cues alone are insufficient wang2021towards . However, existing most vision-language trackers still predominantly adopt local search strategies, restricting the search region to a small neighborhood around the previous target location. While efficient, this localized assumption makes the tracker vulnerable to drift: once the target moves rapidly, undergoes abrupt viewpoint changes, or becomes temporarily occluded, the tracker may fail to rediscover it.
@@
Recent advancements in Multimodal Large Language Models (MLLMs) have opened new possibilities for rethinking the tracking pipeline. MLLMs exhibit remarkable capabilities in global scene understanding, cross-modal reasoning bai2025qwen3 . These strengths suggest that MLLMs could fundamentally reshape long-term tracking, allowing the model to perform global search over the entire image rather than relying on local heuristics. In this paper, we present the first global vision–language tracking framework powered by Multimodal Large Language Models (MLLMs), termed VPTracker. Leveraging the strong semantic reasoning capabilities of MLLMs, our approach can localize the target across the entire image space. Despite their impressive semantic understanding, we observe that large-scale language models inherently lack temporal modeling and spatial continuity awareness—two key elements required for reliable visual tracking. To overcome this limitation, we introduce a location-aware visual prompting mechanism that injects spatial priors directly into the MLLM’s inference process. Specifically, we construct a region-level prompt based on the target’s previous location, guiding the model to focus on the most probable search area while still maintaining the ability to perform global reasoning when necessary. This design preserves the robustness of global search while substantially reducing false positives caused by visually or semantically similar distractors.
@@
The experimental results on the commonly used visual language tracking datasets TNL2K wang2021towards , and TNLLT wang2025reasoningtrack demonstrate that our method has achieved significant improvements in robustness, target re-identification, and long-term stability. More importantly, VPTracker introduces a new perspective of integrating MLLM into vision-language tracking tasks, indicating that future tracking will not only be driven by appearance similarity but also by advanced cross-modal reasoning. In summary, the main contributions of this work are threefold:
@@
  • Extensive experiments on TNL2K and TNLLT demonstrate that our approach achieves superior robustness, improved recognition capability, and strong long-term stability, establishing a new direction for integrating MLLMs into vision–language tracking.
2 Related Work
2.1 Object Tracking
Visual object tracking (VOT) has been extensively studied over the past decade, with research evolving from correlation filter–based methods to modern deep learning–based frameworks ye2022ostrack . Early trackers such as KCF henriques2014high and ECO danelljan2017eco model target appearance through handcrafted features and efficient frequency-domain operations. With the success of deep convolutional networks, Siamese-based trackers (e.g., SiamFC bertinetto2016fully , SiamRPN li2018high , SiamMask hu2023siammask ) became dominant due to their strong template-matching capability and real-time performance. Transformer-based trackers (e.g., TransT chen2021TransT , STARK tang2022SFTransT , OSTrack ye2022ostrack ) further improve robustness by modeling long-range dependencies with self-attention mechanisms.
@@ 2.1 Object Tracking
2.2 Multimodal Large Language Models
Multimodal Large Language Models (MLLMs), such as GPT-4V achiam2023gpt , LLaVA li2024llava , Qwen-VLbai2025qwen3 , and InternVL wang2025internvl3 , have exhibited strong capabilities in high-level scene understanding, cross-modal reasoning, and open-vocabulary recognition wang2025unlocking . Unlike traditional vision-language models that output fixed embeddings, MLLMs are able to perform explicit reasoning, generate spatial descriptions, and interpret complex relationships between objects. These properties make them promising for challenging video tasks such as grounding, captioning, and instruction-following.
@@ 3.1 Problem Formulation
3.2 Global Tracking via MLLMs
To address the aforementioned issues, this paper proposes a global tracking method based on MLLM. Leveraging MLLM’s powerful reasoning capabilities, it searches for targets across the entire video frame given a visual template and corresponding linguistic description, and provides bounding boxes for the detected targets. As illustrated in Fig. 1, our approach relies exclusively on a single multimodal large language model, without requiring any additional components. The prompt employed in our method is presented in Fig. 2.
@@ 3.2 Global Tracking via MLLMs
3.3 Visual Prompt based on Location-aware
@@
4 Experiment
4.1 Implementation Details
@@
4.2 Comparison with State-of-the-art Methods
@@
The TNLLT dataset is a recently proposed benchmark for long-term vision–language tracking, designed to assess a model’s capability to localize target objects over extended video sequences using natural language descriptions. Its test set consists of 50 long and diverse video sequences. To better assess robustness, TNLLT further provides 15 challenging attributes, such as occlusion, deformation, motion blur, illumination variation, fast motion, low resolution, out-of-view, and background clutter. These rich annotations make TNLLT a comprehensive benchmark for evaluating long-term temporal modeling, and cross-modal reasoning in vision–language tracking. We compare VPTracker with existing state-of-the-art trackers in Tab.1. Compared to the state-of-the-art ReasoningTrack, which uses MLLM to update the linguistic description of the object during tracking, we achieve significant improvements over prior methods.
@@
Table 1: Overall tracking performance on TNLLT dataset. The best results are highlighted in bold.
@@
OSTrack ye2022ostrack ECCV22 $B$ 52.1 57.3 63.6
MixFormer cui2022mixformer CVPR22 $B$ 56.0 61.1 67.8
AiATrack gao2022aiatrack ECCV22 $B$ 53.0 56.9 63.5
JointNLT zhou2023joint CVPR23 NL 45.0 47.9 55.5
JointNLT zhou2023joint CVPR23 BL 45.4 48.2 55.9
All-in-one zhang2023all MM23 BL 49.7 54.8 60.6
MMTrack zheng2023toward TCSVT23 BL 55.8 61.8 67.7
CiteTrack citetracker ICCV23 $B$ 50.3 55.3 60.8
ROMTrack cai2023robust ICCV23 $B$ 49.0 54.8 59.9
GRM gao2023generalized CVPR23 $B$ 50.0 55.4 61.3
ODTrack zheng2024odtrack AAAI24 $B$ 54.4 61.3 65.9
EVPTrack shi2024evptrack AAAI24 $B$ 54.4 61.3 65.8
UVLTrack ma2024unifying AAAI24 $B$ 53.1 58.1 64.1
AQATrack xie2024autoregressive CVPR24 $B$ 55.1 61.5 66.5
UVLTrack ma2024unifying AAAI24 NL 51.9 57.4 63.1
UVLTrack ma2024unifying AAAI24 BL 54.4 60.3 66.2
SUTrack sutrack AAAI25 BL 50.8 55.4 60.0
LMTrack xu2025less AAAI25 $B$ 48.7 53.0 59.8
CTVLT feng2025enhancing ICASSP25 BL 60.9 69.5 73.6
@@
The TNL2K dataset is a large dataset used for training and evaluating Vision-Language tracking. It contains 2000 video sequences and 663 words, of which 1300 videos are used for training and 700 videos are used for testing. It introduces two new challenges: adversarial samples and modality switches. We compare VPTracker with existing advanced trackers in Tab.2. we achieve significant improvements over prior methods.
Table 2: Overall tracking performance on TNL2K dataset. The best results are highlighted in bold.
@@
TrDiMP wang2021transformer CVPR21 $B$ - - -
TransT tang2022SFTransT CVPR21 $B$ 50.7 51.7 57.1
AutoMatch zhang2021learn AAAI21 $B$ 47.2 43.5 -
SNLT feng2021siamese CVPR21 BL 27.6 41.9 -
VLTTT guo2022divert NIPS22 BL 53.1 53.3 -
CTRNLT li2022cross CVPR22 BL 44.0 45.0 52.0
JointNLT zhou2023joint CVPR23 BL 56.9 58.1 73.6
OneTrack hong2024onetracker CVPR24 BL 58.0 59.1 -
QueryNLT shao2024context CVPR24 BL 57.8 58.7 75.6
MAVLT shi2025mamba TCSVT25 BL 63.1 66.7 -
ProVLT 10947500 TCSVT25 BL 59.8 61.8 -
TemTrack xie2025robust AAAI25 $B$ 58.8 - -
@@ 4.3 Qualitative Comparison
4.4 Ablation study
To validate the effectiveness of our proposed location-aware visual prompting mechanism, we conduct comparisons on the TNL2K dataset with and without the visual prompt. As shown in Tab. 3, incorporating the visual prompting mechanism leads to consistent and notable improvements across metrics such as AUC, PR, and NPR. To more clearly demonstrate the effectiveness of our proposed VP mechanism, we present a visual comparison of the tracking results in Fig. 4. Since the textual description of the target may only be applicable to the first frame, global tracking without the position-aware visual prompt can lead to incorrect prediction when visually similar objects appear in the scene.
Table 3: Overall tracking performance on TNL2K dataset.
@@ 5 Conclusion
Acknowledgments
This work was partially supported by the National Natural Science Foundation of China under Grant 62072185.